OpenStudy (anonymous):

Ben is a safety inspector for a family gym. Recently, guests have been complaining about the diving board saying that it is leaning towards the water. Ben needs to know whether this diving board is standing up straight or not. Ben measures the height of the diving board to be 15 feet. The pool is 13 feet across and he measures the diagonal from the other side of the pool to the top of the diving board to be 21 feet. Is the diving board at a 90 degree angle with the ground? Explain how you know.

OpenStudy (anonymous):

Please help me..

OpenStudy (anonymous):

 Hi sele-peace, in your problem you can start with the assumption that the diving board is standing up straight. If that's the case, it will make a 90 deg angle with the pool, obviously. If that's true, Pythagorean Law applies, which states that A^2 + B^2 = C^2, where A is the size of the board, B is the size of the pool and C is the size of the diagonal. Note that this only holds for a straight angle (which is what you would like to find out) of the board to the pool. Just fill out the numbers: 15^2+13^2 = 394. If the board would be straight up, the square root of 394 should be 21. As the square root of 394 turns out to be only 19.8 meters, you can conclude that the board is NOT at this 90 degree angle but at a bigger angle. Hope this answers your question ?